es P1-8. Reading a Balance Sheet Southwest Airlines is a U.S. passenger airline focused on do- mestic, point-to-point, low-fare, and short- to medium-haul routes. Its balance sheets as of December 31, 2014 and December 31, 2013 appear in the exhibit below as the basis for the following questions. (Note that the company reports in millions of U.S. dollars.) Required: a. What is Southwest Airlines largest asset? What is its book value as of December 31, 2014? b. Calculate the current ratio e as of December 31, 2014 and December 31, 2013. Why might this ratio be helpful? c. What might be the explanation for the companys air traffic liability? d. What is Southwest Airlines largest long-term liability as of December 31, 2014? e. Calculate the total liabilities as of December 31, 2014. f. Was Southwest Airlines profitable during 2014? How can you tell? 1
media%2F02d%2F02da2a57-6356-4d91-93e7-5a
Answer #1

a)largest asset is flight equipment costs and its book value as of end of 2014 is 18473mn
b))Current ratio 2014=current assets/current liabilites
=4404/5923=0.74
2013=4456/5676=0.79
we can see current ratio decreases in 2014 compared to 2013
c)Air traffic liability is unearned revenue type where the airlines collected money in advance for which the service is not yet provided
d)Deferred income tax is largest longterm liability
e)total liabilites as of 2014= current liability+deferred income tax+construction oblig+other noncurrent lib
=5923+2434+3259+554+1255=13425mn
f)Yes it is profitable as we can retained earnings in 2014 is higher than 2013